3xy-4x+y
x equals -3 and y equals 2
both the 4x and y are squared
: 3xy-4x+y
x equals -3 and y equals 2
both the 4x and y are squared

Answer by jim_thompson5910(9376)


3*x*y-4*x^2+y^2 Start with the given expression.


3*(-3)*2-4*(-3)^2+(2)^2 Plug in x=-3 and y=2


-9*2-4*(-3)^2+(2)^2 Multiply 3 and -3 to get -9.


-18-4*(-3)^2+(2)^2 Multiply -9 and 2 to get -18.


-18-4*9+(2)^2 Raise -3 to the 2nd power to get 9.


-18-36+(2)^2 Multiply 4 and 9 to get 36.


-54+(2)^2 Subtract 36 from -18 to get -54.


-54+4 Raise 2 to the 2nd power to get 4.


-50 Add -54 and 4 to get -50.


So 3*(-3)*2-4*(-3)^2+(2)^2=-50


So 3*x*y-4*x^2+y^2=-50 when x=-3 and y=2